# High-Layer Attention Pruning with Rescaling

### v1

@@ Abstract
Pruning is a highly effective approach for compressing large language models (LLMs), significantly reducing inference latency. However, conventional training-free structured pruning methods often employ a heuristic metric that indiscriminately removes some attention heads across all pruning layers, without considering their positions within the network architecture. In this work, we propose a novel pruning algorithm that strategically prunes attention heads in the model’s higher layers. Since the removal of attention heads can alter the magnitude of token representations, we introduce an adaptive rescaling parameter that calibrates the representation scale post-pruning to counteract this effect. We conduct comprehensive experiments on a wide range of LLMs, including LLaMA3.1-8B, Mistral-7B-v0.3, Qwen2-7B, and Gemma2-9B. Our evaluation includes both generation and discriminative tasks across 27 datasets. The results consistently demonstrate that our method outperforms existing structured pruning methods. This improvement is particularly notable in generation tasks, where our approach significantly outperforms existing baselines.
@@ 1 Introduction
Large language models (LLMs) [78, 67, 42, 88, 26], pre-trained on extensive text data, have achieved surprising performance in downstream tasks such as information retrieval [4], code generation [28], and mathematical reasoning [80, 90, 37]. These LLMs, however, contain a huge number of parameters, resulting in substantially slower inference speed compared to their smaller counterparts.
Many efforts have leveraged pruning algorithms [22, 40, 83, 5, 87, 39, 97, 21, 92, 93, 100] to reduce redundant parameters and lower inference latency without significantly compromising the task performance. Among these, training-free structured pruning methods [61, 2, 74] estimate the importance of parameters using predefined pruning metrics, allowing them to remove less important components without retraining. While most pruned parameters typically come from feedforward networks, these methods also remove some low-importance attention heads. This is especially beneficial for long-context generation tasks, where the attention mechanism becomes the main computational bottleneck due to its quadratic time complexity with respect to sequence length.
However, current pruning algorithms generally prune attention heads across all targeted layers, either by uniformly removing the same heads from each layer [61] or by adaptively pruning more heads in certain layers and fewer in others [2]. These approaches do not consider which attention head positions are actually more redundant and thus better suited for pruning. We argue that the pruning metrics used in these methods are not accurate enough to reliably identify the most redundant attention heads. In this work, we aim to address the following question:
@@
We approach this problem by analyzing the relative importance of different layers in LLMs. Recent studies [62, 101, 27, 31, 72] demonstrate that pruning blocks or entire layers in the high layers of LLMs often yields better performance than pruning those in the low layers. Building on these findings, we propose a training-free structured pruning algorithm, called HARP (High-layer Attention Rescaled Pruning), that removes the attention heads, specifically the query and key parameters, in the higher layers. After pruning, the attention mechanism in these layers can be skipped entirely. However, bypassing the attention computation may alter the output magnitude, as the value representations are no longer combined through weighted averaging based on attention scores over the full sequence. To mitigate this effect, we introduce a rescaling parameter in the attention residual block to calibrate the token representation magnitude after pruning.
We evaluate the effectiveness of our approach by applying it to LLaMA3.1-8B, Mistral-7B-v0.3, Qwen2-7B, and Gemma2-9B. Extensive experiments on 27 representative downstream tasks show that our method outperforms popular training-free structured pruning algorithms, including Magnitude Pruning [30], LLM-Pruner [61], FLAP [2], and Wanda [74]. As illustrated on the left side of Figure 1, our method, HARP, significantly outperforms these baselines on generation tasks under the same attention head pruning ratio. More notably, the results on the right side of Figure 1 show that applying rescaling significantly improves performance compared to the unscaled variant, demonstrating the importance of rescaling for properly adjusting the representation magnitude.
@@ 2.1 Pruning Attention Heads for Long-Context Modeling
When $N<d_{\mathrm{ff}}$, the FFN’s $\mathcal{O}\left(Ndd_{\mathrm{ff}}\right)$ complexity becomes the main time bottleneck. Conversely, when $N>d_{\mathrm{ff}}$, the $\mathcal{O}\left(Nd^{2}\right)$ complexity of the self-attention module is dominant. In long-context tasks, prompt lengths can reach up to 2 million tokens [44], making the operations in the self-attention module extremely time-expensive. By pruning attention heads [8], we can achieve significant improvements in computational efficiency for long-context tasks while only pruning a small fraction of the model’s parameters.
@@ 2.2 High-Layer Attention Heads Are Less Important
We begin with the following theoretical analysis.
@@
Define the average pairwise cosine similarity among tokens in a matrix $\mathbf{H}\in\mathbb{R}^{N\times d}$ as
$\operatorname{Sim}(\mathbf{H}):=\frac{2}{N(N-1)}\sum_{1\leq i<j\leq N}\frac{\mathbf{h}_{i}^{\top}\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|\|\mathbf{h}_{j}\|}.$ (1)
This measures the global similarity of token representations.
@@
As $\operatorname{Sim}(\mathbf{H}^{(\ell)})\to 1$ for a high-layer $\ell$ in a Transformer, the attention scores in that layer tend to uniform distributions, reducing their discriminative capacity. Therefore, the attention mechanism in higher layers becomes less informative than in lower layers where $\operatorname{Sim}(\mathbf{H}^{(\ell)})<1$.
Proof can be found in Appendix A. This theorem suggests that as token similarities increase and attention scores approach uniform distributions, attention heads become less effective at capturing accurate dependencies between tokens.
@@
Recent studies [62, 101, 27, 31, 72, 57, 95, 41, 14, 46] have introduced various heuristic metrics to evaluate the importance of blocks within each layer of LLMs. A common finding across these works is that the parameters in lower layers tend to be more important than those in higher layers, lending empirical support to our theoretical analysis. One such metric is the Block Importance (BI) score [62]:
@@
The BI score typically decreases with increasing layer index $\ell$, indicating the reduced impact of higher-layer parameters. In this work, we build upon these analysis to build our pruning algorithm.
@@ 2.3 Methodology
In this section, we introduce HARP, a novel structured pruning algorithm specifically designed to remove attention heads in the higher layers of LLMs. Our approach focuses on pruning only the query and key parameters in these layers. As a result, we can bypass the self-attention mechanism for these pruned layers. The computation within the self-attention module can be formulated as:
$\mathbf{H}=\mathbf{H}\mathbf{W}_{V}\mathbf{W}_{O}.$ (3)
It is important to note that many LLMs adopt grouped-query attention (GQA) [1]. To accommodate such an attention mechanism, a slight modification is necessary: we replicate the value matrix $\mathbf{V}=\mathbf{H}\mathbf{W}_{V}$ a total of $g-1$ times, where $g$ denotes the number of query groups.
@@
The aggregation step in self-attention, which computes a weighted average of value representations across the sequence, can change the overall magnitude [48], typically measured by the Frobenius norm of the hidden states $\mathbf{H}$.
@@
Let $\mathbf{H}\in\mathbb{R}^{N\times d}$ denote the token representations, and let $\hat{\mathbf{A}}\in\mathbb{R}^{N\times N}$ be a row-normalized attention matrix (i.e., each row sums to 1). Consider the aggregation step in self-attention defined by $\mathbf{H}^{\prime}=\hat{\mathbf{A}}\mathbf{H}$. Then,
@@
with equality if and only if each row of $\hat{\mathbf{A}}$ is a one-hot vector. This result indicates that the aggregation operation typically changes the magnitude of token representations.
Based on this proposition, we argue that the magnitude of token representations can change after pruning, since the aggregation step is bypassed. Inspired by normalization approaches [38, 6], which maintain the magnitude of token representations within a stable range to prevent gradient explosion or vanishing during pretraining, we similarly adjust the magnitude after pruning. This adjustment keeps the token representation magnitude within a stable range relative to their pre-pruning state, helping preserve model performance during inference. We introduce a rescaling parameter $\alpha$ into the residual block of the self-attention module as follows:
$\mathbf{H}=\mathbf{H}+\alpha\mathbf{H}\mathbf{W}_{V}\mathbf{W}_{O}.$ (5)
@@
1:  Input: Base LLM, number of layers to prune $P$, search values $\alpha\in\{0,0.1,\ldots,1.0\}$
2:  Initialize: $\bm{\alpha}_{\text{best}}\in\mathbb{R}^{P}=[1.0,1.0,\ldots,1.0]$
3:  $L=$ total number of layers in the model
4:  for $\ell=L-1$ down to $L-P$ do
5:     $\mathrm{PPL}_{\text{best}}=\infty$
6:     for each $\alpha\in\{0,0.1,\ldots,1.0\}$ do
7:        Set rescaling parameter $\bm{\alpha}_{\text{best}}[L-1-\ell]=\alpha$ for layer $\ell$
8:        $\mathrm{PPL}=\operatorname{Perplexity}(\text{LLM},\bm{\alpha}_{\text{best}})$
9:        if $\mathrm{PPL}<\mathrm{PPL}_{\text{best}}$ then
10:           $\mathrm{PPL}_{\text{best}}=\mathrm{PPL}$
11:           $\bm{\alpha}_{\text{best}}[L-1-\ell]=\alpha$
12:        end if
13:     end for
14:  end for
15:  Output: Optimal rescaling parameters $\bm{\alpha}_{\text{best}}\in\mathbb{R}^{P}$ for each pruned layer
Determining the optimal value of $\alpha$ is challenging, as traditional retraining methods are not applicable due to the non-differentiable nature of $\alpha$. While recent works have employed Zeroth-Order Optimization [29, 96] to estimate gradients during fine-tuning, we propose a simpler and more efficient solution: a greedy search strategy. Our method selects the optimal $\alpha$ by evaluating the perplexity of the pruned LLM. To simplify the search process, we adopt a top-down approach: starting from the topmost layer, we determine the optimal $\alpha$ for that layer and fix it before proceeding to search for $\alpha$ in the next lower layer. This layer-by-layer greedy strategy forms a sequential for-loop that significantly reduces the overall search space. The entire procedure is implemented as a grid search, as illustrated in Algorithm 1. Developing more advanced search strategies remains an open direction for future work.
@@ 3 Experiments
In this section, we present a comprehensive evaluation of our proposed pruning algorithm, evaluating its performance across four LLMs and eleven diverse benchmarks. By conducting experiments on various LLMs and benchmarks, we aim to establish consistent and reliable results.
@@ 3.1 Experimental Setup
We evaluate HARP against four baselines: Magnitude pruning [30], LLM-Pruner [61], FLAP [2], and Wanda [74]. We build upon the FLAP to extend Magnitude Pruning and Wanda to structured pruning, referring to these variants as Mag-sp and Wanda-sp. Mag-SP, LLM-Pruner, and Wanda-SP prune attention heads uniformly across all targeted layers by removing the same number of heads from each layer, whereas FLAP adaptively prunes more heads in some layers and fewer in others.
Table 1: Performance on generation tasks across LLaMA3.1-8B, Mistral-7B-v0.3, Qwen2-7B, and Gemma2-9B (%).
@@
Table 2: Performance on discriminative tasks across LLaMA3.1-8B, Mistral-7B-v0.3, Qwen2-7B, and Gemma2-9B (%).
@@
We conduct experiments on four GQA-based LLMs: LLaMA3.1-8B [78], Mistral-7B-v0.3 [42], Qwen2-7B [88], and Gemma2-9B [26]. For LLaMA3.1-8B and Mistral-7B-v0.3, we set the attention head pruning ratio to 1/4, whereas for Qwen2-7B and Gemma2-9B, it is set to 1/7. Therefore, we use HARP to prune 8 layers for LLaMA3.1-8B, 4 layers for Qwen2-7B, 8 layers for Mistral-7B-v0.3, and 6 layers for Gemma2-9B. We find that our method achieves a smaller pruning ratio compared to the baselines. To address this and ensure a fair comparison, we integrate FLAP with our HARP, leveraging FLAP to prune FFN parameters in our implementation. Aside from this, FFN parameters are not pruned in the baseline methods.
@@
It is important to note that the official implementation [^1] https://github.com/CASIA-IVA-Lab/FLAP does not support GQA-based LLMs. Therefore, we use a modified implementation [^2] https://github.com/nyunAI/FLAP for our experiments. Furthermore, the official implementation [^3] https://github.com/horseee/LLM-Pruner of LLM-Pruner is available only for LLaMA3.1-8B. Our HARP is compatible with any attention mechanism.
@@
For our evaluation, we employ the widely-used lm-evaluation-harness [^4] https://github.com/EleutherAI/lm-evaluation-harness package [24] to conduct experiments on both generation and discriminative tasks. Our generation tasks include 5-shot TriviaQA [45], 5-shot NaturalQuestions [50], 5-shot GSM8K [17], 4-shot MATH-hard [34, 52], 3-shot COT [81], and BBH [76]. For discriminative tasks, we use 5-shot WinoGrande [69], 25-shot ARC-Challenge [16], 0-shot BoolQ [15], 0-shot OpenBookQA [64], 0-shot PIQA [9], and 5-shot MMLU [33]. We report the accuracy for these tasks as recommended by the lm-evaluation-harness package. We also evaluate long-context generation performance on the LongBench [^5] https://github.com/THUDM/LongBench/tree/main benchmark [7], which includes 16 datasets. Details about benchmarks can be found in Appendix C.

### v2

@@ Abstract
Pruning is a highly effective approach for compressing large language models (LLMs), significantly reducing inference latency. However, conventional training-free structured pruning methods often employ a heuristic metric that indiscriminately removes some attention heads across all pruning layers, without considering their positions within the network architecture. In this work, we propose a novel pruning algorithm that strategically prunes attention heads in the model’s higher layers. Since the removal of attention heads can alter the magnitude of token representations, we introduce an adaptive rescaling parameter that calibrates the representation scale post-pruning to counteract this effect. We conduct comprehensive experiments on a wide range of LLMs, including LLaMA3.1-8B, Mistral-7B-v0.3, Qwen2-7B, and Gemma2-9B. Our evaluation includes both generation and discriminative tasks across 27 datasets. The results consistently demonstrate that our method outperforms existing structured pruning methods. This improvement is particularly notable in generation tasks, where our approach significantly outperforms existing baselines. Code is available at https://github.com/SongtaoLiu0823/HARP.
@@ 1 Introduction
Large language models (LLMs) (Touvron et al., 2023; OpenAI et al., 2023; Jiang et al., 2023; Yang et al., 2024; Gemma2-Team et al., 2024), pre-trained on extensive text data, have achieved surprising performance in downstream tasks such as information retrieval (Asai et al., 2024), code generation (Guo et al., 2024), and mathematical reasoning (Wang et al., 2023; Yang et al., 2023b; Huang et al., 2024). These LLMs, however, contain a huge number of parameters, resulting in substantially slower inference speed compared to their smaller counterparts.
Many efforts have leveraged pruning algorithms (Frantar and Alistarh, 2023; Jaiswal et al., 2023; Xia et al., 2024; Ashkboos et al., 2024; Xu et al., 2024; Jaiswal et al., 2024a; Zhang et al., 2024b; Dong et al., 2024b; Yin et al., 2024a; b; Zhao et al., 2024) to reduce redundant parameters and lower inference latency without significantly compromising the performance. Among these, training-free structured pruning methods (Ma et al., 2023; An et al., 2024; Sun et al., 2024b) estimate the importance of parameters using predefined pruning metrics, allowing them to remove less important components without retraining. While most pruned parameters typically come from feedforward networks, these methods also remove some low-importance attention heads. This is especially beneficial for long-context tasks, where the attention mechanism becomes the main computational bottleneck due to its quadratic time complexity with respect to sequence length.
However, current pruning algorithms generally prune attention heads across all targeted layers, either by uniformly removing the same heads from each layer (Ma et al., 2023) or by adaptively pruning more heads in certain layers and fewer in others (An et al., 2024). These approaches do not consider which attention head positions are actually more redundant and thus better suited for pruning. We argue that the pruning metrics used in these methods are not accurate enough to reliably identify the most redundant attention heads. In this work, we aim to address the following question:
@@
We approach this problem by analyzing the relative importance of different layers in LLMs. Recent studies (Men et al., 2024; Zhong et al., 2024; Gromov et al., 2024; He et al., 2024; Siddiqui et al., 2024) demonstrate that pruning blocks or entire layers in the high layers of LLMs often yields better performance than pruning those in the low layers. Building on these findings, we propose a training-free structured pruning algorithm, called HARP (High-layer Attention Rescaled Pruning), that removes the attention heads, specifically the query and key parameters, in the higher layers. After pruning, the attention mechanism in these layers can be skipped entirely. However, bypassing the attention computation may alter the output magnitude, as the value representations are no longer combined through weighted averaging based on attention scores over the full sequence. To mitigate this effect, we introduce a rescaling parameter in the attention residual block to calibrate the token representation magnitude after pruning.
We evaluate the effectiveness of our approach by applying it to LLaMA3.1-8B, Mistral-7B-v0.3, Qwen2-7B, and Gemma2-9B. Extensive experiments on 27 representative downstream tasks show that our method outperforms popular training-free structured pruning algorithms, including Magnitude Pruning (Han et al., 2015), LLM-Pruner (Ma et al., 2023), FLAP (An et al., 2024), and Wanda (Sun et al., 2024b). As illustrated on the left side of Figure 1, our method, HARP, significantly outperforms these baselines on generation tasks under the same attention head pruning ratio. More notably, the results on the right side of Figure 1 show that applying rescaling significantly improves performance compared to the unscaled variant, demonstrating the importance of rescaling for properly adjusting the representation magnitude.
@@ 2.1 Pruning Attention Heads for Long-Context Modeling
When $N<d_{\mathrm{ff}}$, the FFN’s $\mathcal{O}\left(Ndd_{\mathrm{ff}}\right)$ complexity becomes the main time bottleneck. Conversely, when $N>d_{\mathrm{ff}}$, the $\mathcal{O}\left(N^{2}d\right)$ complexity of the self-attention mechanism dominates. In long-context tasks, prompt lengths can reach up to 2 million tokens (Jin et al., 2025), making the operations in the self-attention module extremely time-expensive. By pruning attention heads (Bansal et al., 2022), we can achieve significant improvements in computational efficiency for long-context tasks while only pruning a small fraction of the model’s parameters.
@@ 2.2 High-Layer Attention Heads Are Less Important
In this subsection, we formalize what we mean by “tokens becoming similar” in a layer and how this affects the usefulness of self-attention. Let $\mathbf{H}^{(\ell)}\in\mathbb{R}^{N\times d}$ denote the hidden representation matrix at layer $\ell$, whose $i$-th row $\mathbf{h}_{i}^{(\ell)}$ corresponds to the representation of the $i$-th token. Throughout this paper, when we say that “the token representations are (almost) the same”, we specifically mean that their pairwise cosine similarities are close to $1$, i.e., all token vectors are nearly parallel in the representation space.
@@
The average pairwise cosine similarity among tokens in a matrix $\mathbf{H}\in\mathbb{R}^{N\times d}$ is defined as
$\operatorname{Sim}(\mathbf{H}):=\frac{2}{N(N-1)}\sum_{1\leq i<j\leq N}\frac{\mathbf{h}_{i}^{\top}\mathbf{h}_{j}}{\|\mathbf{h}_{i}\|_{2}\|\mathbf{h}_{j}\|_{2}}.$ (1)
A value $\operatorname{Sim}(\mathbf{H})\approx 1$ indicates that all token representations are almost colinear (up to scaling), while a smaller value means that the representations point in more diverse directions. This quantity therefore measures the global level of similarity among token representations.
@@
When $\operatorname{Sim}(\mathbf{H}^{(\ell)})\to 1$, token representations in layer $\ell$ become nearly parallel. Since queries and keys are linear projections of $\mathbf{H}^{(\ell)}$, their dot products across positions become almost constant, so the row-wise softmax yields nearly uniform attention weights. The head thus degenerates into averaging value vectors and cannot distinguish tokens. Conversely, if $\operatorname{Sim}(\mathbf{H}^{(\ell)})$ is bounded away from $1$, attention remains non-uniform and can model meaningful token-to-token dependencies.
A formal proof of Theorem 1 is provided in Appendix A. Intuitively, as the hidden states within a layer become more and more aligned (large $\operatorname{Sim}(\mathbf{H}^{(\ell)})$), the attention operation behaves like a simple averaging operator that mixes already similar representations. This phenomenon is analogous to over-smoothing in deep GNNs, where repeated averaging causes node embeddings to collapse into a low-dimensional subspace. As a result, attention heads in very high layers, where representations are highly aligned, are less informative than those in lower layers where representations remain more diverse.
@@
Recent studies (Men et al., 2024; Zhong et al., 2024; Gromov et al., 2024; He et al., 2024; Siddiqui et al., 2024; Liu et al., 2024b; Zhang et al., 2024a; Jaiswal et al., 2024b; Chen et al., 2024b; Kim et al., 2024a) have introduced various heuristic metrics to evaluate the importance of blocks within each layer of LLMs. A common finding across these works is that the parameters in lower layers tend to be more important than those in higher layers, lending empirical support to our theoretical analysis. One such metric is the Block Importance (BI) score (Men et al., 2024):
@@
The BI score typically decreases with increasing layer index $\ell$, indicating the reduced impact of higher-layer parameters. In this work, we use these empirical observations as additional evidence that pruning higher-layer attention heads tends to be less harmful than pruning lower-layer ones, and we explicitly incorporate this layer-wise asymmetry into our pruning design.
@@ 2.3 Methodology
In this section, we introduce HARP, a novel structured pruning algorithm specifically designed to remove attention heads in the higher layers of LLMs. Our approach focuses on pruning only the query and key parameters in these layers. As a result, we can bypass the self-attention mechanism for these pruned layers. Concretely, when the attention of a layer is fully pruned, the layer no longer computes attention scores or performs aggregation; instead, it only applies the value and output projections. The computation within the self-attention module can then be formulated as $\mathbf{H}=\mathbf{H}\mathbf{W}_{V}\mathbf{W}_{O}.$
@@
The aggregation step in self-attention computes, for each token, a weighted average of value vectors over all positions in the sequence. This operation not only mixes information across tokens but can also change the overall scale of the hidden representations (Kipf and Welling, 2017). We measure this scale using the Frobenius norm of the hidden state matrix $\mathbf{H}$.
@@
Let $\mathbf{H}\in\mathbb{R}^{N\times d}$ denote the token representations, and let $\hat{\mathbf{A}}\in\mathbb{R}^{N\times N}$ be a row-normalized attention matrix (i.e., each row is non-negative and sums to $1$). Consider the aggregation step in self-attention defined by $\mathbf{H}^{\prime}=\hat{\mathbf{A}}\mathbf{H}$. In general, we have
@@
and equality holds if and only if every row of $\hat{\mathbf{A}}$ is a one-hot vector (so that each output token simply copies one input token without averaging). This means that, except for this degenerate case, the aggregation step changes the magnitude of token representations.
@@
Input: Base LLM, number of layers to prune $P$, search values $\alpha\in\{0,0.1,\ldots,1.0\}$
Initialize: $\bm{\alpha}_{\text{best}}\in\mathbb{R}^{P}=[1.0,1.0,\ldots,1.0]$
$L=$ total number of layers in the model
for $\ell=L-1$ down to LPL-P do
  $\mathrm{PPL}_{\text{best}}=\infty$
  for each $\alpha\in\{0,0.1,\ldots,1.0\}$ do
   Set rescaling parameter $\bm{\alpha}_{\text{best}}[L-1-\ell]=\alpha$ for layer $\ell$
   $\mathrm{PPL}=\operatorname{Perplexity}(\text{LLM},\bm{\alpha}_{\text{best}})$
   if $\mathrm{PPL}<\mathrm{PPL}_{\text{best}}$ then
    $\mathrm{PPL}_{\text{best}}=\mathrm{PPL}$
    $\bm{\alpha}_{\text{best}}[L-1-\ell]=\alpha$
   end if
  end for
end for
Output: Optimal rescaling parameters $\bm{\alpha}_{\text{best}}\in\mathbb{R}^{P}$ for each pruned layer
A detailed proof of Proposition 1 is also given in Appendix A (“Proof of Proposition 1”). The proposition implies that removing the aggregation step (by pruning queries and keys and bypassing attention) will in general alter the scale of the hidden states compared to the original model.
Based on this proposition, we argue that the magnitude of token representations can change after pruning, since the aggregation step is bypassed. Inspired by normalization approaches (Ioffe and Szegedy, 2015; Ba et al., 2016), which maintain the magnitude of token representations within a stable range to prevent gradient explosion or vanishing during pretraining, we similarly adjust the magnitude after pruning. This adjustment keeps the token representation magnitude within a stable range relative to their pre-pruning state, helping preserve model performance during inference. We introduce a rescaling parameter $\alpha$ into the residual block of the self-attention module as $\mathbf{H}=\mathbf{H}+\alpha\mathbf{H}\mathbf{W}_{V}\mathbf{W}_{O}.$
Determining the optimal value of $\alpha$ is challenging, as we treat $\alpha$ as a non-differentiable, layer-wise hyperparameter rather than a trainable parameter. Instead of retraining the model, we propose a simple and efficient greedy search strategy that selects $\alpha$ by directly evaluating the perplexity of the pruned LLM on a held-out corpus.
To simplify the search process, we adopt a top-down approach: starting from the topmost layer (layer $L$), we determine the optimal $\alpha$ for that layer while keeping all other layers unpruned or fixed; we then fix this value and proceed to search for $\alpha$ in the next lower layer. This layer-by-layer greedy strategy forms a sequential for-loop that significantly reduces the overall search space, as shown in Algorithm 1.
In practice, we implement this procedure as a grid search over a small set of candidate values $\{0,0.1,\ldots,1.0\}$, which requires only a modest number of forward passes (perplexity evaluations) and is therefore much cheaper than any training-based method. Designing more advanced search strategies (e.g., Bayesian optimization or differentiable relaxations) is an interesting open direction for future work.
@@ 3 Experiments
Table 1: Comparison of FLAP with existing pruning algorithms on LLMs.
Method Weight Update Calibration Data Pruning Metric
Magnitude $|\mathbf{W}_{ij}|$
Wanda $|\mathbf{W}_{ij}|\cdot\|\mathbf{X}_{j}\|_{2}$
FLAP $\frac{1}{N-1}\sum_{n=1}^{N}\left(\mathbf{X}_{n,j,:}^{\ell}-\overline{\mathbf{X}}_{:,j,:}^{\ell}\right)^{2}\cdot\left\|\mathbf{W}_{:,j}^{\ell}\right\|_{2}^{2}$
HARP ppl
@@ 3.1 Experimental Setup
We evaluate HARP against four baselines: Magnitude pruning (Han et al., 2015), LLM-Pruner (Ma et al., 2023), FLAP (An et al., 2024), and Wanda (Sun et al., 2024b), as shown in Table 1. We build upon the FLAP to extend Magnitude Pruning and Wanda to structured pruning, referring to these variants as Mag-sp and Wanda-sp. Mag-SP, LLM-Pruner, and Wanda-SP prune attention heads uniformly across all targeted layers by removing the same number of heads from each layer, whereas FLAP adaptively prunes more heads in some layers and fewer in others.
Table 2: Performance on generation tasks across LLaMA3.1-8B, Mistral-7B-v0.3, Qwen2-7B, and Gemma2-9B (%). Best is bold.
@@
Table 3: Performance on discriminative tasks across LLaMA3.1-8B, Mistral-7B-v0.3, Qwen2-7B, and Gemma2-9B (%). Best is bold
@@
We conduct experiments on four GQA-based LLMs: LLaMA3.1-8B (Touvron et al., 2023), Mistral-7B-v0.3 (Jiang et al., 2023), Qwen2-7B (Yang et al., 2024), and Gemma2-9B (Gemma2-Team et al., 2024). For LLaMA3.1-8B and Mistral-7B-v0.3, we set the attention head pruning ratio to 1/4, whereas for Qwen2-7B and Gemma2-9B, it is set to 1/7. Therefore, we use HARP to prune 8 layers for LLaMA3.1-8B, 4 layers for Qwen2-7B, 8 layers for Mistral-7B-v0.3, and 6 layers for Gemma2-9B. We find that our method achieves a smaller pruning ratio compared to the baselines. To address this and ensure a fair comparison, we integrate FLAP with our HARP, leveraging FLAP to prune FFN parameters in our implementation. Aside from this, FFN parameters are not pruned in the baseline methods.
@@
It is important to note that the FLAP official implementation [^1] https://github.com/CASIA-IVA-Lab/FLAP does not support GQA-based LLMs. Therefore, we use a modified implementation [^2] https://github.com/nyunAI/FLAP for our experiments. Furthermore, the official implementation [^3] https://github.com/horseee/LLM-Pruner of LLM-Pruner is available only for LLaMA3.1-8B. Our HARP is compatible with any attention mechanism.
@@
Cost analysis and comparison with baselines.
Our implementation performs a layer-wise grid search, evaluating $K$ candidate $\alpha$ values per layer using perplexity-based forward passes. Thus, the total cost scales as $L\times K$ forward passes for an $L$-layer model. For the 8-layer setting used in our experiments with K=10K{=}10, this amounts to $8\times 10=80$ forward passes in total. In contrast, inference-only pruning baselines such as Mag-sp, Wanda-sp, and FLAP generally require only a single calibration/inference pass to collect importance statistics, i.e., roughly $\sim 1\times$ the cost of one forward pass. Overall, HARP introduces a one-time search overhead (about $80\times$ more forward passes than single-pass baselines in this 8-layer case), but remains substantially cheaper than training-based methods; moreover, this cost is amortized since the search is performed only once to select a good $\alpha$.
@@
For our evaluation, we employ the widely-used lm-evaluation-harness package (Gao et al., 2024) to conduct experiments on both generation and discriminative tasks. Our generation tasks include 5-shot TriviaQA (Joshi et al., 2017), 5-shot NaturalQuestions (Kwiatkowski et al., 2019), 5-shot GSM8K (Cobbe et al., 2021), 4-shot MATH-hard (Hendrycks et al., 2021b; Lewkowycz et al., 2022), 3-shot COT (Wei et al., 2022a), and BBH (Suzgun et al., 2023). For discriminative tasks, we use 5-shot WinoGrande (Sakaguchi et al., 2021), 25-shot ARC-Challenge (Clark et al., 2018), 0-shot BoolQ (Clark et al., 2019), 0-shot OpenBookQA (Mihaylov et al., 2018), 0-shot PIQA (Bisk et al., 2020), and 5-shot MMLU (Hendrycks et al., 2021a). We report the accuracy for these tasks as recommended by the lm-evaluation-harness package. We also evaluate long-context generation performance on the LongBench benchmark (Bai et al., 2024), which includes 16 datasets. Details about benchmarks can be found in Appendix C.